As investment with a 3year life and a cost of #120000 generates revenue of #25000 in year 1 #45000 in year 2 and #65000 in year 3. If the discount rate 8%

(i) what is the NPV of the investment.

(ii) state whether the investment should be accepted or not and why.

(iii) using 12% as your second discount rate, to solve the remaining part of the question under IRR.


1
Expert's answer
2021-08-02T15:12:02-0400

Solution


i.) PV=120000


YEAR 1= "\frac{25000}{1.08}=23148.14"


YEAR 2="\frac{45000}{1.08^2}=38580.25"


YEAR 3="\frac{65000}{1.08^3}=51599.10"


"NPV= -120000+23148.15+38580.25+51599.10"


NPV1= -6672.5


ii.) The investment should not be accepted because the NPV is a negative.


iii.) Using 12% as the discount rate

PV= 120000


1st"=\frac{25000}{1.12}=22321.42"


2nd"=\frac{45000}{1.12^2}=35873.72"


3rd"=\frac{65000}{1.12^3}=46265.72"


NPV2"=-120000+22321.43+35873.72+46265.72"

= -15539.13


"IRR=R_1+\frac{NPV_1\times(R_2-R_1)}{{NPV_1-NPV_2}}"


IRR"=6.5" %